Lilian9667:

5x-7x+6=-2(x-3) A.) 0 B.) 3/4 C.) 2 D.) infinitely many solutions

Lilian9667:

hmm?

SadHoodie:

smackz:

5x - 7x + 6 = -2(x - 3) Distribute the -2 5x - 7x + 6 = -2x + 6 Combine like terms on the left side -2x + 6 = -2x + 6 Add 2x to both sides 6 = 6 so d

sxdsouls:

1: Simplify both sides of the equation. 5x−7x+6=−2(x−3) 5x+−7x+6=(−2)(x)+(−2)(−3)(Distribute) 5x+−7x+6=−2x+6 (5x+−7x)+(6)=−2x+6(Combine Like Terms) −2x+6=−2x+6 −2x+6=−2x+6 2: Add 2x to both sides. −2x+6+2x=−2x+6+2x 6=6 3: Subtract 6 from both sides. 6−6=6−6 0=0 All real numbers are solutions.

supie:

So the first step would be to distribute \(5x−7x+6=−2(x−3)\) \(5x+−7x+6=(−2)(x)+(−2)(−3)\) \(5x+−7x+6=−2x+6\) Then combine like terms \((5x+−7x)+(6)=−2x+6\) \(−2x+6=−2x+6\) Then take \(2x\) add it to both sides of the equation \(−2x+6+2x=−2x+6+2x\) \(=6=6\) Subtract 6 from both sides of the equation \(6−6=6−6\)

sxdsouls:

No direct answers smh.

sxdsouls:

yh np,

Lilian9667:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls yh np, \(\color{#0cbb34}{\text{End of Quote}}\) Wait I don't see a 6 there

Lilian9667:

There's No 6 there so can you explain that

sxdsouls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Lilian9667 \(\color{#0cbb34}{\text{Originally Posted by}}\) @sxdsouls yh np, \(\color{#0cbb34}{\text{End of Quote}}\) Wait I don't see a 6 there \(\color{#0cbb34}{\text{End of Quote}}\) 2: Add 2x to both sides. −2x+6+2x=−2x+6+2x 6=6 3: Subtract 6 from both sides. 6−6=6−6 0=0

sxdsouls:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @supie So the first step would be to distribute \(5x−7x+6=−2(x−3)\) \(5x+−7x+6=(−2)(x)+(−2)(−3)\) \(5x+−7x+6=−2x+6\) Then combine like terms \((5x+−7x)+(6)=−2x+6\) \(−2x+6=−2x+6\) Then take \(2x\) add it to both sides of the equation \(−2x+6+2x=−2x+6+2x\) \(=6=6\) Subtract 6 from both sides of the equation \(6−6=6−6\) \(\color{#0cbb34}{\text{End of Quote}}\) most likely what he said i guess

Lilian9667:

ok

sxdsouls:

got it ?

Lilian9667:

yup

sxdsouls:

mk